Calculate the mass percent of different elements in sodium sulphate (Na2SO4).

Grade:12

1 Answers

Gaurav
9 years ago
Molar mass of the whole compound = 142g
Mass percentage of Na =\frac{Mass\ of\ two\ sodium\ atom}{Molar\ mass\ of\ compound}\times 100
= (2 x 23/142) x 100 = 32.4

Similarly,
Mass percentage of S = (32/142)x100 = 22.54
Mass percentage of O = 100 – 32.4 – 24.1